OpenStudy (anonymous):

What is the maximum volume of an open rectangular box (with no top face) if its surface area is 1 square foot?

OpenStudy (anonymous):

Since there is a constraint im pretty sure you have to use Lagrange Multipliers

OpenStudy (anonymous):

V(l,w,h) = lwh Vl(l,w,h) = hw Vh(l,w,h) = lw Vw(l,w,h) = lh

OpenStudy (anonymous):

l = length, w = width, h = height

OpenStudy (anonymous):

so the constraint would be S(l,w,h) = 2(lh + wh) + wl since the base is only counted once

OpenStudy (anonymous):

S being the surface area and V the volume

OpenStudy (anonymous):

Sl(l,w,h) = 2h + w Sh(l,w,h) = 2l + 2w S(l,w,h) = 2h + l

OpenStudy (anonymous):

so the equations would look like this hw = (lamda)(2h + w) lw = (landa)(2l + 2w) lh = (lamda(2h + l) 2(lh + hw) + wl = 1

OpenStudy (anonymous):

i dont know how to quite go from here to find l, w, and h

OpenStudy (anonymous):

Any ideas?

OpenStudy (threshsupport):

so one panel is 1 m^2

OpenStudy (anonymous):

panel?

OpenStudy (anonymous):

it's a box that is missing the top face so it's open

OpenStudy (threshsupport):

ok

OpenStudy (threshsupport):

i cubic meter

OpenStudy (threshsupport):

one

OpenStudy (anonymous):

i think multiply eq (1) with l and (2) with h to get a relation... multiply eq(1) with l and (3) with w to get a relation... and use those in equation (4) See if that helps

OpenStudy (anonymous):

when you say multiply do you mean like solve eq 1 for w and plug into eq 2?

OpenStudy (anonymous):

lwh = (lambda)*l*(2h + w) lwh = (lambda)*h*(2l + 2w) solve these two :)

OpenStudy (anonymous):

hmmm ok let me try that, i hadn't thought of doing it that way

OpenStudy (anonymous):

sure...

OpenStudy (anonymous):

ok got l = 2h from that

OpenStudy (anonymous):

so basically i just do the same with 2 and 3 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok thank you and is there a general way to know when to do a trick like that?

OpenStudy (anonymous):

or is it just a matter of doing a bunch of those types of problems and figuring it out?

OpenStudy (anonymous):

i think if you practice a lot you would figure it out

OpenStudy (anonymous):

ok thank you sir